Calibration status for Run 15

Speaker : Gene V. Buren ( BNL )


Talk time : 12:35, Duration : 00:15
 Calibration status for Run 15

_________________
  • pp200
    • TPC
      • Alignment & T0: not expected to be different from Run 14 [done]
        • Will be checked using field-on cosmic ray data for later improvements to calibration
      • Drift velocities: issues caused by weak laser signals during the run were resolved modified track-seed finding for laser files [done]
        • Might modify the automatic calibration for Run 16
      • SpaceCharge & GridLeak: [in progress]
        • Highest luminosities delivered to-date for a pp200 Run; asymmetric due to HFT support materials
        • Four potentially different run setups:
          production_pp200long_2015, production_pp200long2_2015, production_pp200long3_2015, production_pptrans_2015
          • long2 and trans are the major datasets
          • long was early on, and is being treated here with lower priority
          • long3 is just a few runs (ignoring)
        • Early stages of the calibration compared long2 and trans and found the calibration looks quite similar between the two, so a single calibration will be done
        • A final iteration of the calibration is in-progress right now (~1 day away) using more restrictive cuts against pile-up
      • dE/dx: needs a production touching nearly all runs [not done; awaits BeamLine]
      • hit errors: [not done; awaits SC&GL]
    • BeamLine
      • Preliminary in place, calibrated during Run [done]
        • Verified to be not far off during TPC SC&GL calibration
      • Awaits TPC SC&GL [not done]
      • A better, HFT-constrained BeamLine should be calibrated after HFT alignment [not done]
    • HFT
      • Status tables need a production touching nearly all runs [not done - no feedback on plans at this time]
        • Could overlap with TPC dE/dx, but may need a supplement to span time across the runs (i.e. first event to last event) as was done for Run 14 (supplement doesn't need tracking)
      • Alignment: required as PXL was removed for maintenance and reinserted
        • Internal alignment using zero field cosmic rays [in progress]
        • HFT-to-TPC alignment using field-on cosmic rays [not done - no feedback on plans at this time]
    • TOF
      • BTOF timing will use Run 14 calibration [done]
      • VPD [not done]
      • BTOF-to-TPC alignment: needs a calibration production which can overlap with TPC dE/dx [not done]
    • EMCs
      • Pedestals and status tables [in progress]
  • pAu200 & pAl200 [identical statuses]
    • TPC
      • Drift velocities [done]
      • SpaceCharge & GridLeak [not done]
      • dE/dx [not done]
    • BeamLine
      • Same status as pp200 [not done]
    • HFT
      • Status tables: possibly do at the same time as pp200 [not done]
    • TOF
      • VPD [not done]
      • BTOF can use same calibrations as pp200 [not done]
    • EMC
      • Pedestals and status tables [not done]
  • Fixed target (AuAu)
    • Database considerations needed
      • Fixed target runs were only over a short period of time in the middle of the pAu running
      • pAu database values need to be picked up both before and after the fixed target runs (need copied)
    • TPC
      • Drift velocities [done]
      • SpaceCharge & GridLeak not expected to be relevant
      • dE/dx [done]
    • Beamline
      • Fixed target wants a modified vertex finder [not done]
    • HFT
      • Won't be used; not relevant
    • TOF
      • VPD calibration check needed [in progress]
      • BTOF need not be separately calibrated
    • EMC
      • Pedestals and status tables [not done]

-Gene